# Time Series Anomaly Detection in the Frequency Domain with Statistical Reliability

### v1

@@ Acknowledgments
This work was partially supported by MEXT KAKENHI (20H00601), JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
@@ Appendix A Determination of Penalty Parameters in the Optimization Problem
In this section, we derive the penalty term in (5) used for detecting CP candidates in the frequency domain. We begin by assuming that the true mean vector of $\bm{f}^{(d)}$ for frequency $d\in\{0,\dots,D-1\}$ is piecewise constant as follows
@@
where $\mu_{\text{seg}}^{(d)}\left(\tau_{k-1}^{(d)}:\tau_{k}^{(d)}\right)$ represents the true mean of the $k$-th segment in frequency $d$, and $\sigma_{f}^{2}$ denotes the known variance such that $\sigma_{f}^{2}=M\sigma^{2}$ by the property of DFT777 Although we impose the assumption of the piecewise constant mean structure for model selection based on BIC, the theoretical validity of $p$-values obtained by our proposed SI method is guaranteed even when this assumption does not hold. . Then, we introduce BIC for deriving the values of the penalty parameters $\bm{\beta}$ in (5). Given the assumptions of (21), the unknown parameters $\bm{\theta}$ of the sequence $\bm{f}^{(d)}$ are expressed as
@@
$\gamma=\kappa M\sigma^{2}\log T,$
where $\kappa$ is a hyper-parameter determined by the user based on problem settings. We heuristically set $\kappa=0.5$ in the synthetic data experiments, and $\kappa=3$ in the real data experiment.
@@ B.1 Proof of Theorem 1
Proof. According to the second condition in (16), we have
@@
$\Leftrightarrow\bm{X}$ $=\mathcal{U}(\bm{x})+\mathcal{V}(\bm{X})z,$
$\Leftrightarrow\bm{X}$ $=\mathcal{U}(\bm{x})+\mathcal{V}(\bm{x})z,\,(\because\mathcal{V}(\bm{X})=\mathcal{V}(\bm{x}))$
@@
where $\bm{a}=\mathcal{U}(\bm{x}),\bm{b}=\mathcal{V}(\bm{x})$, and $z=T_{k}(\bm{X})=\sigma^{-1}||P_{k}\bm{X}||$.
Then, we have
$\mathcal{X}$ =$=\{\bm{X}\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\}$
=$=\{\bm{X}\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\bm{X}=\bm{a}+\bm{b}z,z\in\mathbb{R}\}$
=$=\{\bm{X}=\bm{a}+\bm{b}z\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{a}+\bm{b}z)=\mathcal{A}(\bm{x}),z\in\mathbb{R}\}$
=$=\{\bm{X}=\bm{a}+\bm{b}z\in\mathbb{R}^{N}\,|\,z\in\mathcal{Z}\}.$
Therefore, we obtain the result in Theorem 1.
@@ B.2 Proof of Theorem 2
Proof. The sampling distribution of the test statistic conditional on $\mathcal{A}(\bm{X})$ and $\mathcal{Q}(\bm{X})$ denoted by
@@
follows a truncated $\chi$-distribution with $\mathrm{tr}(P_{k})$ degrees of freedom and the truncation region $\mathcal{Z}$ defined in (19). Thus, by applying the probability integral transform, under the null hypothesis,
@@
$\mathbb{P}_{\rm{H}_{0},k}\left(p_{k}^{\rm{selective}}\leq\alpha\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\right)=\alpha,\,\forall\alpha\in[0,1].$
Next, for any $\alpha\in[0,1]$, we have
@@ Appendix C Identification of Truncation Region
In general, it is difficult to identify the truncation region $\mathcal{Z}$ in (19) directly because conditioning only on the result of the CP candidate selection, i.e., $\mathcal{A}(\bm{a}+\bm{b}z)=\mathcal{A}(\bm{x})$, is intractable. In this case, we have to enumerate all patterns where $\mathcal{A}(\bm{x})$ appears as a result of simulated annealing algorithm, which is computationally impractical. To address this issue, we first compute the region conditioned on the process of the algorithm $\mathcal{A}$. That is to say, it is guaranteed that the process remains identical within the region. This additional conditioning is often denoted as “over-conditioning” because it is redundant for valid inference. In the case of SI with over-conditioning, the type I error rate can still be controlled at the significance level, while the power tends to be low (Lee et al., 2016; Liu et al., 2018; Duy & Takeuchi, 2022). Therefore, we apply an efficient line search method based on parametric programming to compute $\mathcal{Z}$ where the redundant conditioning is removed for the purpose of improving the power (Duy & Takeuchi, 2022). In the following, we first compute the over-conditioned region in Appendix C.1, and then identify the truncation region $\mathcal{Z}$ using parametric programming in Appendix C.2.
C.1 Characterizing using Over-conditioning
Since we only need to consider one-dimensional data space in $\mathbb{R}^{N}$ as indicated in (17), we define the over-conditioned region $\mathcal{Z}^{\text{oc}}$ where the process of the algorithm $\mathcal{A}$ remains unchanged as
@@ C.1 Characterizing using Over-conditioning
Over-conditioning on dynamic programming. We compute $\{r\in\mathbb{R}\,|\,\mathcal{S}_{\text{DP}}(r)=\mathcal{S}_{\text{DP}}(z)\}$ by conditioning on all the operations based on the Bellman equation which is used in the dynamic programming algorithm to obtain the optimal solution in (6). Since the Bellman equation consists of the cost $\mathcal{C}(\cdot)$ and the penalty $\bm{\beta}$, the condition is finally represented by a quadratic inequality, as in the following discussion of simulated annealing. The detail derivation is presented in Duy et al. (2020).
Over-conditioning on simulated annealing containing the preliminary experiment. Considering the algorithm of simulated annealing, we find that the procedure depends on $\bm{X}=\bm{a}+\bm{b}z$ at only one specific point, that is, the Metropolis algorithm, which is given by
status $\text{status is}\begin{cases}\text{Acceptance}&\left(\exp\left(\frac{-\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)}{c_{i}}\right)>\theta_{i,l}\right)\\ \vspace{-1mm}\\ \text{Rejection}&\left(\exp\left(\frac{-\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)}{c_{i}}\right)\leq\theta_{i,l}\right)\end{cases},$
where $\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)=E(\bm{\mathcal{T}}_{i,l},\bm{a}+\bm{b}r)-E(\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)$, and $\bm{\mathcal{T}}_{i,l-1}$, $\bm{\mathcal{T}}_{i,l}$ respectively represent CP candidates before and after a transition to the neighborhood in the $l$-th iteration with a parameter $\theta_{i,l}$ for the $i$-th temperature $c_{i}$. Note that while the transitions are accepted when either $\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)\leq 0$ or $\exp(-\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)/c_{i})>\theta_{i,l}$, the latter case includes the former. Thus, to compute the region $\{r\in\mathbb{R}\,|\,\mathcal{S}_{\text{SA}}(r)=\mathcal{S}_{\text{SA}}(z)\}$ where the process of simulated annealing is identical, we need to condition on all results of the Metropolis algorithm at each temperature. This condition consists of multiple inequalities as follows
@@
$=\bigcap_{i=0}^{I_{z}}\bigcap_{l=1}^{L_{z}}\left\{r\in\mathbb{R}\,\Bigg{|}\,\begin{cases}\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\theta_{i,l})<0&\text{(status is Acceptance)}\\ \Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\theta_{i,l})\geq 0&\text{(status is Rejection)}\end{cases}\right\},$ (24)
@@
$C^{(d)}_{s+1:e}=c_{\text{sym}}^{(d)}\sum_{t=s+1}^{e}\bm{u}^{(d)}_{s+1:e,t}{\bm{u}_{s+1:e,t}^{{\ast^{(d)}}^{\top}}}\in\mathbb{R}^{N\times N},$
@@
and $\bm{u}_{s+1:e,t}^{\ast^{(d)}}$ is the complex conjugate of $\bm{u}^{(d)}_{s+1:e,t}$. Therefore, the objective function $E(\bm{\mathcal{T}},\bm{a}+\bm{b}r)$ in (5) can be rewritten as follows
@@
$=\bigcap_{i=0}^{I_{z}^{+}}\bigcap_{l=1}^{L_{z}^{+}}\left\{r\in\mathbb{R}\,\Bigg{|}\,\begin{cases}\Delta E(\bm{\mathcal{T}}_{i,l}^{+},\bm{\mathcal{T}}_{i,l-1}^{+},\bm{a}+\bm{b}r)+c_{i}^{+}\ln(\theta_{i,l}^{+})<0&\text{(status is Acceptance)}\\ \Delta E(\bm{\mathcal{T}}_{i,l}^{+},\bm{\mathcal{T}}_{i,l-1}^{+},\bm{a}+\bm{b}r)+c_{i}^{+}\ln(\theta_{i,l}^{+})\geq 0&\text{(status is Rejection)}\end{cases}\right\},$
where $\bm{\mathcal{T}}_{i,l}^{+},c_{i}^{+},\theta_{i,l}^{+},I_{z}^{+}$ and $L_{z}^{+}$ in the preliminary experiment correspond to the respective parameters in (24), and $\bm{\mathcal{T}}_{i,0}^{+}=\bm{\mathcal{T}}^{\text{init}}$.
@@ C.2 Parametric Programming
Having derived $\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)$ in the previous analysis, the region $\mathcal{Z}$ in (19) conditioned on the result of the algorithm $\mathcal{A}$ can be obtained using a computational method called parametric programming as follows
@@
The overall procedure for computing selective $p$-values of the detected CP candidate locations is presented in Algorithm 3. Furthermore, the line search method based on (26) for obtaining the region $\mathcal{Z}$ required for the computation of $p_{k}^{\text{selective}}$ is detailed in Algorithm 4. An overview of the proposed search method is shown in Figure 10.
Algorithm 3 SI for detected CP candidate locations
1:$\bm{x}$
2:$\bm{\mathcal{T}}^{\text{det}}\leftarrow\mathcal{A}(\bm{x})$ in (8)
3:Obtain $\bm{\tau}^{\text{det}}$ by (4)
4:for $\tau_{k}^{\text{det}}\in\bm{\tau}^{\text{det}}$ do
5:   $\mathcal{Z}\leftarrow$ compute_solution_path$\left(\bm{x},\bm{\mathcal{T}}^{\text{det}},\tau_{k}^{\text{det}}\right)$
6:   Compute $p_{k}^{\text{selective}}$ by (20)
7:end for
8:$\left\{\left(\tau_{k}^{\text{det}},p_{k}^{\text{selective}}\right)\right\}_{k=1}^{K}$
@@
1:$\bm{x},\bm{\mathcal{T}}^{\text{det}}$ and $\tau_{k}^{\text{det}}$
@@
13:$\mathcal{Z}$
@@
Appendix D Details of the Numerical Experiments

### v2

@@ Acknowledgments
This work was partially supported by JST CREST (JPMJCR21D3, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), and RIKEN Center for Advanced Intelligence Project.
@@ Appendix A Determination of Penalty Parameters in the Optimization Problem
In this section, we derive the penalty term in (5) used for detecting CP candidates in the frequency domain. We begin by assuming that the true mean vector of $\bm{f}^{(d)}$ for frequency $d\in\{0,\dots,D-1\}$ is piecewise constant as follows:
@@
where $\mu_{\text{seg}}^{(d)}\left(\tau_{k-1}^{(d)}:\tau_{k}^{(d)}\right)$ represents the true mean of the $k$-th segment in frequency $d$, and $\sigma_{f}^{2}$ denotes the known variance such that $\sigma_{f}^{2}=M\sigma^{2}$ by the property of DFT666 Although we impose the assumption of the piecewise constant mean structure for model selection based on BIC, the theoretical validity of $p$-values obtained by our proposed SI method is guaranteed even when this assumption does not hold. . Then, we introduce BIC for deriving the values of the penalty parameters $\bm{\beta}$ in (5). Given the assumptions of (18), the unknown parameters $\bm{\theta}$ of the sequence $\bm{f}^{(d)}$ are expressed as
@@
$\gamma=\kappa M\sigma^{2}\log T,$ (20)
where $\kappa$ is a hyper-parameter determined by users based on problem settings. We heuristically set $\kappa=0.5$ in the synthetic data experiments, and $\kappa=3$ in the real data experiment.
@@ B.1 Proof of Theorem 1
Proof.
According to the condition on $\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})$, i.e., $\mathcal{U}(\bm{X})=\mathcal{U}(\bm{x})$ and $\mathcal{V}(\bm{X})=\mathcal{V}(\bm{x})$, we have
@@
$\Leftrightarrow\bm{X}$ $=\mathcal{U}(\bm{x})+\mathcal{V}(\bm{X})z$
$\Leftrightarrow\bm{X}$ $=\mathcal{U}(\bm{x})+\mathcal{V}(\bm{x})z~{}~{}~{}(\because\mathcal{V}(\bm{X})=\mathcal{V}(\bm{x}))$
@@
where $\bm{a}=\mathcal{U}(\bm{x}),\bm{b}=\mathcal{V}(\bm{x})$, and $z=T_{k}(\bm{X})=\sigma^{-1}||P_{k}\bm{X}||$. Then, we have
$\{\bm{X}\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\}$
$=$ $\{\bm{X}\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\bm{X}=\bm{a}+\bm{b}z,z\in\mathbb{R}\}$
$=$ $\{\bm{X}=\bm{a}+\bm{b}z\in\mathbb{R}^{N}\,|\,\mathcal{A}(\bm{a}+\bm{b}z)=\mathcal{A}(\bm{x}),z\in\mathbb{R}\}$
$=$ $\{\bm{X}=\bm{a}+\bm{b}z\in\mathbb{R}^{N}\,|\,z\in\mathcal{Z}\}.$
Therefore, by noting that $\|P_{k}\bm{s}\|$ is zero, we obtain
$T(\bm{X})\mid\{\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\}\sim\mathrm{TC}(\mathrm{tr}(P_{k}),\mathcal{Z}).$
@@ B.2 Proof of Theorem 2
Proof.
The sampling distribution of the test statistic conditional on $\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x})$ and $\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})$ denoted by
@@
is a truncated $\chi$-distribution with the degrees of freedom $\mathrm{tr}(P_{k})$ and the truncation region $\mathcal{Z}$ defined in Theorem 1. Thus, by applying the probability integral transform, under the null hypothesis,
@@
$\mathbb{P}_{\rm{H}_{0},k}\left(p_{k}^{\rm{selective}}\leq\alpha\,|\,\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x}),\mathcal{Q}(\bm{X})=\mathcal{Q}(\bm{x})\right)=\alpha,\,\forall\alpha\in(0,1).$
Next, for any $\alpha\in(0,1)$, we have
@@ Appendix C Identification of Truncation Region
In general, it is difficult to identify the truncation region $\mathcal{Z}$ in Theorem 1 directly because conditioning only on the result of the CP candidate selection, i.e., $\mathcal{A}(\bm{a}+\bm{b}z)=\mathcal{A}(\bm{x})$, is intractable. In this case, we have to enumerate all patterns where $\mathcal{A}(\bm{x})$ appears as a result of simulated annealing algorithm, which is computationally impractical. To address this issue, we first compute the region conditioned on the process of the algorithm $\mathcal{A}$. That is to say, it is guaranteed that the process remains identical within the region. This additional conditioning is often denoted as “over-conditioning” because it is redundant for valid inference. In the case of SI with over-conditioning, the type I error rate can still be controlled at the significance level, while the power tends to be low (Lee et al., 2016; Liu et al., 2018; Duy & Takeuchi, 2022). Therefore, we apply an efficient line search method based on parametric programming to compute $\mathcal{Z}$ where the redundant conditioning is removed for the purpose of improving the power (Duy & Takeuchi, 2022). In the following, we first compute the over-conditioned region in Appendix C.1, and then identify the truncation region $\mathcal{Z}$ using parametric programming in Appendix C.2.
C.1 Over-Conditioning
Since we only need to consider one-dimensional data space in $\mathbb{R}^{N}$, we define the over-conditioned region $\mathcal{Z}^{\text{oc}}$ where the process of the algorithm $\mathcal{A}$ remains unchanged as
@@ C.1 Over-Conditioning
Over-conditioning on dynamic programming.
We compute $\{r\in\mathbb{R}\,|\,\mathcal{S}_{\text{DP}}(r)=\mathcal{S}_{\text{DP}}(z)\}$ by conditioning on all the operations based on the Bellman equation which is used in the dynamic programming algorithm to obtain the optimal solution in (7). Since the Bellman equation consists of the cost $\mathcal{C}(\cdot)$ and the penalty $\bm{\beta}$, the condition is finally represented by a quadratic inequality, as in the following discussion of simulated annealing. The detail derivation is presented in Duy et al. (2020).
Over-conditioning on simulated annealing containing the preliminary experiment.
Considering the algorithm of simulated annealing, we find that the procedure depends on $\bm{X}=\bm{a}+\bm{b}z$ at only one specific point, that is, the Metropolis algorithm, which is given by
status $\text{{status} is}\begin{cases}\text{Acceptance}&\text{if $\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\xi_{i,l})<0$},\\ \vspace{-1mm}\\ \text{Rejection}&\text{if $\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\xi_{i,l})\geq 0$},\end{cases}$
where $\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)=E(\bm{\mathcal{T}}_{i,l},\bm{a}+\bm{b}r)-E(\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)$, and $\bm{\mathcal{T}}_{i,l-1}$, $\bm{\mathcal{T}}_{i,l}$ respectively represent CP candidates before and after a transition to the neighborhood in the $l$-th iteration with a parameter $\xi_{i,l}$ for the $i$-th temperature $c_{i}$. Thus, to compute the region $\{r\in\mathbb{R}\,|\,\mathcal{S}_{\text{SA}}(r)=\mathcal{S}_{\text{SA}}(z)\}$ where the process of simulated annealing is identical, we need to condition on all results of the Metropolis algorithm at each temperature. This condition consists of multiple inequalities as follows:
@@
$=\bigcap_{i=0}^{I_{z}}\bigcap_{l=1}^{L_{z}}\left\{r\in\mathbb{R}\,\Bigg{|}\,\begin{cases}\Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\xi_{i,l})<0&\text{{if {status} is Acceptance}}\\ \Delta E(\bm{\mathcal{T}}_{i,l},\bm{\mathcal{T}}_{i,l-1},\bm{a}+\bm{b}r)+c_{i}\ln(\xi_{i,l})\geq 0&\text{{if {status} is Rejection}}\end{cases}\right\},$ (22)
@@
$C^{(d)}_{s+1:e}=c_{\text{sym}}^{(d)}\sum_{t=s+1}^{e}\bm{u}^{(d)}_{s+1:e,t}{\bm{u}_{s+1:e,t}^{{{(d)}\ast}\top}}\in\mathbb{R}^{N\times N},$
@@
and $\bm{u}_{s+1:e,t}^{(d)\ast}$ is the complex conjugate of $\bm{u}^{(d)}_{s+1:e,t}$. Therefore, the objective function $E(\bm{\mathcal{T}},\bm{a}+\bm{b}r)$ in (5) can be rewritten as follows:
@@
$=\bigcap_{i=0}^{I_{z}^{+}}\bigcap_{l=1}^{L_{z}^{+}}\left\{r\in\mathbb{R}\,\Bigg{|}\,\begin{cases}\Delta E(\bm{\mathcal{T}}_{i,l}^{+},\bm{\mathcal{T}}_{i,l-1}^{+},\bm{a}+\bm{b}r)+c_{i}^{+}\ln(\xi_{i,l}^{+})<0&\text{{if {status} is Acceptance}}\\ \Delta E(\bm{\mathcal{T}}_{i,l}^{+},\bm{\mathcal{T}}_{i,l-1}^{+},\bm{a}+\bm{b}r)+c_{i}^{+}\ln(\xi_{i,l}^{+})\geq 0&\text{{if {status} is Rejection}}\end{cases}\right\},$
where $\bm{\mathcal{T}}_{i,l}^{+},c_{i}^{+},\xi_{i,l}^{+},I_{z}^{+}$, and $L_{z}^{+}$ in the preliminary experiment correspond to the respective parameters in (22), and $\bm{\mathcal{T}}_{i,0}^{+}=\bm{\mathcal{T}}^{\text{init}}$.
@@ C.2 Parametric Programming
Having derived $\mathcal{Z}^{\text{oc}}(\bm{a}+\bm{b}z)$ in the previous analysis, the region $\mathcal{Z}$ in Theorem 1 conditioned on the result of the algorithm $\mathcal{A}$ can be obtained using a computational method called parametric programming as follows:
@@
The line search method based on (24) for obtaining the region $\mathcal{Z}$ required for the computation of $p_{k}^{\text{selective}}$ is detailed in Algorithm 3. Furthermore, the overall procedure for computing selective $p$-values of the detected CP candidate locations is presented in Algorithm 4. An overview of the proposed search method is shown in Figure 9.
@@
1:Time seaquence $\bm{x}$, CP candidates $\bm{\mathcal{T}}$, CP candidate location $\tau_{k}$
@@
13:Truncation region $\mathcal{Z}$
Algorithm 4 SI for Detected CP Candidate Locations
1:Time seaquence $\bm{x}$
2:$\bm{\mathcal{T}}\leftarrow\mathcal{A}(\bm{x})$
3:Obtain $\bm{\tau}$ by (4)
4:for $\tau_{k}\in\bm{\tau}$ do
5:   $\mathcal{Z}\leftarrow$ compute_solution_path$\left(\bm{x},\bm{\mathcal{T}},\tau_{k}\right)$
6:   Compute $p_{k}^{\text{selective}}$ by (17)
7:end for
8:Detected CP candidate locations and the corresponding selective $p$-values $\left\{\left(\tau_{k},p_{k}^{\text{selective}}\right)\right\}_{k=1}^{K}$
@@
Appendix D Details of Numerical Experiments